OpenStudy (anonymous):

Perform the indicated operation. (36m^2 - 60m + 25) ÷ (6m - 5) = _____

OpenStudy (anonymous):

answers 6m + 5 6m - 5 R. 50 6m + 15 R. 100 6m - 5

OpenStudy (anonymous):

6m - 5 because the numerator is a perfect square of 6m - 5

OpenStudy (callisto):

(36m^2 - 60m + 25) ÷ (6m - 5) = [ (6m)^2 - 2(6m)(5) + 5^2 ] ÷ (6m - 5) = (6m-5)^2 ÷ (6m - 5) = (6m-5)^(2-1) = ... ?